In year 1 (the current year), LAA Inc. made a charitable donation of $147,000 to the American Red Cross (a qualifying charity). For the year, LAA reported taxable income of $638,000, which included a $147,000 charitable contribution deduction (before limitation), a $66,000 dividends received deduction, and a $16,300 net operating loss carryover from year 0.

What is LAA Inc.’s charitable contribution deduction for year 1?
Answer #1

LAA's charitable contribution deduction = $85100

Maximum charitable contribution deduction = 10% of taxable income before charitable contribution and dividends received deduction. However, NOL carryovers are deducted from taxable income to determine charitable contribution deduction.

LAA's modified taxable income = 638000+147000+66000 = $851000

LAA's charitable contribution deduction =851000*10% = $85100 ($700,000 × 10%).

The remaining $61900 (147000-85100) is carried over for up to five years.